OpenStudy (anonymous):

What is the 22nd term of the arithmetic sequence where a1 = 8 and a9 = 56?

hartnn (hartnn):

what have you tried? know the required formulas?

OpenStudy (anonymous):

i am confused about what it wants i have a list of formulas next to me

hartnn (hartnn):

"What is the 22nd term" so you will use n'th term formula "of the arithmetic sequence" of the arithmetic sequence

hartnn (hartnn):

\(a_n = a_1 +(n-1)d\)

hartnn (hartnn):

we are given a9 so plug in n= 9 in that formula

OpenStudy (anonymous):

sooo: a of n = a1+(9-1)d

hartnn (hartnn):

n= 9 so, left side = a9 = 56 and a1 is also given, right!

OpenStudy (anonymous):

a1 is 8 right?

hartnn (hartnn):

yep

hartnn (hartnn):

56 = 8+8d

OpenStudy (anonymous):

so d is 6

hartnn (hartnn):

yep, now use the same formula again to find 22d term = a22 just plug in n=22

hartnn (hartnn):

22nd*

OpenStudy (anonymous):

an=8+(22−1)6

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

so 134??

hartnn (hartnn):

correct :)